OpenStudy (anonymous):

Working alone, Ryan can dig a 10ft. by 10ft. hole in five hours. Castel can dig the same hole in six hours. How long would it take them if they worked together?

OpenStudy (chelseatheruler):

of hours working together—x: x/5 + x/6 = 1 6x + 5x = 30 11x = 30 x = 30/11 or 2 8/11 Answer: 2 8/11 hrs OR 2 hrs 43 min 38 2/11 sec

OpenStudy (anonymous):

thanks so much! I suck on work problems